# ASTERIX: Module for modelling the water flow on vegetated hillslopes

### v1

@@ Abstract
The paper presents ASTERIX, an open source software for numerical integration of an extended Saint-Venant system of equations used as a mathematical tool to model the water flow from laboratory up to large-scale spatial domains applying physically-based principles of fluid mechanics. Many in-situ observations have shown that the plant cover plays a key roll in controlling the hydrological flux at a catchment scale. The plant roots facilitate the infiltration processes, the canopy intercept some proportion of rain, and plant stems slow down the flow. In case of heavy rains, the infiltration and interception processes cease in a short time, the remaining rainfall gives rise to the Hortonian overland flow and the flash flood is thus initiated. In this context, the following problem is also addressed in the article: how do the gradient of soil surface and the density of the plant cover influence the water dynamics in the Hortonian flow? The mathematical model and ASTERIX were kept as simple as possible in order to be accessible to a wide range of stakeholders interested in understanding the complex processes behind the water flow on hillslopes covered by plants.
The software is written in C programming language and it is free under GNU license. It was tested on a series of benchmark problems, laboratory experiments, and theoretical problems; and the results have shown a good agreement with the theoretical or measured data.
@@
Software Required: 32 or 64 bit Linux operating system, gcc compiler, SDL libraries.
@@
Availability: free downloadable program sources at
http://www.ima.ro/software/ASTERIX_flow_2D.tar.xz
@@ 1 Introduction
Given all that, it is mandatory for all of us to understand the hydrological cycles to proper manage the water resources in order to protect our life and material well-being.
Mathematical modelling the hydrological processes is an effective way to estimate the risks regarding these water-related natural hazards. By modelling, one can make scenarios, can evaluate the level of the risk exposure, and can provide important information regarding an economical water use (e.g. to agronomists).
@@
In addition, there are many software products that implement mathematical models, some of them are commercial and some others are free. From the last category, we can mention [9, 10] for hydrology-related R and Python tools and resources. Among the most known models in the class of physical models, we mention SHE [11], MIKE-SHE [12], KINEROS [13], VIC [14], PRMS [15], SWASHES [16].
@@
The novelty of adding a pointwise distributed porosity function to model the presence of plant cover allows the study of water dynamics for various scenarios with different vegetation distributions. Such a function can be also used to model the building distribution in urban flood studies [17, 18].
@@
The numerical scheme of the model behind this software was subjected to several internal and external validation tests. Tests such as steady flow over a bump, the oscillating solution of Thacker problem, and the solution to the Riemann problem of the shallow-water equations were used for the internal validation which consists in comparing the numerical results with the exact solutions of the model. For the external validation, we performed tests to compare the numerical results with laboratory measured data [20, 21]: dam break flow over a bump and in a L-shaped channel, downslope flow through rigid vegetation. A good agreement between the numerical and the exact/measured data was obtained in all these cases. A qualitative analysis emphasizing the similarity between the flow evolution on vegetated soil given by our numerical model and the behavior of the observed phenomenon was also performed to strengthen to confidence in the model.
@@
This paper is organized as follows: in Section 2 we present the PDE (Partial Differential Equation) system of the Saint-Venant model with vegetation and the numerical scheme built to approximate this model and implemented in ASTERIX. The software is described in Section 3 and some numerical applications add new examples for additional validation on the method and the software used in Section 4. The last section presents final remarks and conclusions.
@@ 2.1 Mathematical model
The Saint-Venant model with porosity we consider in this paper is based on mass and momentum balance equations and takes into account the variation of the soil surface, the plant cover, and the rain as water source. It reads as:
$\begin{split}\partial_{t}(\theta h)+\partial_{1}(\theta hv_{1})+\partial_{2}(\theta hv_{2})&=\mathfrak{M},\\ \partial_{t}(\theta hv_{a})+\partial_{1}(\theta hv_{a}v_{1})+\partial_{2}(\theta hv_{a}v_{2})+\theta h\partial_{a}{w}&=\mathfrak{t}^{p}_{a}+\mathfrak{t}^{s}_{a},\quad a=1,2.\end{split}$ (1)
@@ 2.2 Numerical scheme
The numerical scheme introduced in [19] to approximate the solution of model (1) casts into the general class of method of lines. We firstly construct a discretization of the spatial variables and differential operators only, and leave the time variable continuous. Then, we define a time discretization scheme to integrate the ODE (Ordinary Differential Equation) system previously obtained. A first-order FVM (Finite Volume Method) is used for the spatial discretization and a fractional step method for the time discretization.
@@
$\begin{array}[]{l}(v_{a})_{(i,j)}:=\displaystyle\frac{v_{a\,i}+v_{a\,j}}{2},\quad a=1,2,\\ (v_{n})_{(i,j)}:=\boldsymbol{v}_{(i,j)}\cdot\boldsymbol{n}_{(i,j)},\vspace{1mm}\\ w_{(i,j)}:=\displaystyle\frac{w_{i}+w_{j}}{2},\end{array}$ (14)
@@
where $l_{(i,j)}$ is the length of the common interface $(i,j)$ between the cells $\omega_{i}$ and $\omega_{j}$. Combining the spatial discretization with the time fractional step method gives the full discrete scheme associated to (1) of the form
$\begin{split}\sigma_{i}(\theta_{i}h_{i})^{n+1}&=\sigma_{i}(\theta_{i}h_{i})^{n}+\triangle t_{n}{\cal L}_{i}((h,\boldsymbol{v})^{n})+\sigma_{i}\triangle t_{n}\mathfrak{M}_{i}(t^{n+1},h^{n+1}),\\ \sigma_{i}(\theta_{i}h_{i}v_{a\,i})^{n+1}&=\sigma_{i}(\theta_{i}h_{i}v_{a\,i})^{n}+\triangle t_{n}\left({\cal J}_{a\,i}((h,\boldsymbol{v})^{n})+{\cal S}_{a\,i}((h,w)^{n})\right)-\triangle t_{n}\sigma_{i}{\cal K}_{i}(h^{n+1})|\boldsymbol{v}_{i}^{n+1}|v^{n+1}_{a\,i},\end{split}$ (16)
for $a=1,2$, where a generic $\psi_{i}$ represents the value of $\psi$ on $\omega_{i}$ for $i\in\overline{1,N}$, $\varphi^{n}$ is the value of $\varphi$ at the moment of time $t^{n}$, $\triangle t_{n}:=t^{n+1}-t^{n}$ is the time step, ||$\cdot$|| ⋅ | is the euclidean norm, and
@@
To advance a time step, one must solve the system (16) for each cell $\omega_{i}$, which basically reduces to determining the water depth $h_{i}^{n+1}$ and velocity $\boldsymbol{v}_{i}^{n+1}$ at the time $t^{n+1}$ based on the known values $h_{i}^{n}$, $\boldsymbol{v}_{i}^{n}$, $t^{n}$ and $\triangle t_{n}$. This is accomplished in three steps:
  1. 1.
    Set $\triangle t_{n}$ subject to the restriction
    $\triangle t_{n}<\tau_{n},$
    where
    $\tau_{n}=CFL\displaystyle\frac{\phi_{\rm min}}{c^{n}_{\rm max}},$ (18)
    with
    $c_{i}=|\boldsymbol{v}|_{i}+\sqrt{gh_{i}},\quad c_{\rm max}=\max\limits_{i}\{c_{i}\},\quad\phi_{\rm min}=\min\limits_{i}\left\{\displaystyle\frac{\sigma_{i}}{\sum\limits_{j\in{\cal N}(i)}l_{(i,j)}}\right\}.$ (19)
  2. 2.
    Calculate $h_{i}^{n+1}$ using the mass balance relation of (16).
  3. 3.
The time step limitation is imposed by the hyperbolic character of the shallow-water equations and by the positivity-preserving requirement of the water depth. $CFL$ is a number between 00 and $1$ (the Courant-Friedrichs-Lewy condition). The reader is referred to A for extended details on the last two steps.
When the software is used to study the Riemann Problem or models that develop shock type discontinuity, it is recommended to use an augmented linear moment flux ${\cal J}^{\boldsymbol{v}}_{a\,i}$ instead of ${\cal J}_{a\,i}(h,\boldsymbol{v})$:
This update is based on the artificial viscosity
where
$c_{(i,j)}=\max\{c_{i},c_{j}\},\quad\mu_{(i,j)}=\displaystyle\frac{2\theta_{i}h_{i}\theta_{j}h_{j}}{\theta_{i}h_{i}+\theta_{j}h_{j}}.$ (22)
@@ 3 Software description
ASTERIX is a software developed in C on a Linux operating system, being dedicated to the simulation of water flow on vegetated slopes. It is based on the previously described numerical scheme for the mathematical model (1) and uses a regular hexagonal network for the spatial discretization. SDL libraries are required for visualizing the time distribution of water layer depth or velocities. ASTERIX is distributed under GPL free software license.
The source code of ASTERIX is a collection of files, most of them containing many specific functions. These component functions can be logically grouped into three categories: Data In, Running and Data Out. One can improve the performance or add new actions by modifying the existent functions or adding new ones. We believe the source code and the accompanying User Guide contain sufficient details to allow a C advised programmer to make the desired modifications.
@@
Domain geometry and the cell partition of the flow domain. The topographic information of the soil provided by a GIS (Geographic Information System) raster is ported to a regular hexagonal raster along this step. One edits the file hexa_parameters.dat in order to establish the number of hexagonal cells on the first row of the hexagonal grid. The number of lines of the hexagonal raster is then automatically calculated depending on the geometry of the GIS raster.
@@
Boundary conditions. The software allows various boundary conditions (free discharge, given mass flux, impermeable border, Dirichlet, and combinations of these) and contains some examples in this sense. The boundary conditions are themselves a rather difficult challenge due to the hyperbolic type of the system of equations. Furthermore, identifying the parts of the border on which different conditions must be established represents another difficulty.
@@
Table 1: Folder structure of ASTERIX with the key files.
@@
Encoded names in ASTERIX Vars &
params
Observations
Unknowns water_vector[$\cdot$] $h(t,\boldsymbol{x})$ Must be given at $t=0$.
@@
External data elevation_vector[$\cdot$] $z(\boldsymbol{x})$ Quantities defining the terrain configuration.
theta_veg[$\cdot$] $\theta(\boldsymbol{x})$
alpha_terr $\alpha_{s}$
@@
4 Numerical examples
A suite of examples is presented in [19] for the purpose of validating the numerical scheme. In what follows, we add some new examples for additional testing on the method, as well as on the software used.
4.1 Flow on radial symmetric surfaces
@@
The numerical solution was obtained with ASTERIX on a hexagonal raster with approximately $900000$ cells of radius $0.115$ m. A comparison between the analytic and numerical water depth and velocity distributions of the stationary solution of the water flow along a radial section is presented in Fig. 4. For the errors between the two solutions we used the formulas
$\epsilon_{h}=\frac{1}{N}\sum_{i=1}^{N}{\left|\frac{h_{n}(r_{i})-h_{a}(r_{i})}{h_{a}(r_{i})}\right|},\qquad\epsilon_{v}=\frac{1}{N}\sum_{i=1}^{N}{\left|\frac{v_{n}(r_{i})-v_{a}(r_{i})}{v_{a}(r_{i})}\right|},$ (26)
Crater type soil Hillock type soil
$\epsilon_{h}$ $0.0055$ $0.0111$
$\epsilon_{v}$ $0.0040$ $0.0018$
@@
4.2 Thacker Problem
The next 2D experiment, Thacker Problem, is a powerful test to check the ability of our numerical scheme to catch and highlight the propagation of a wet/dry front. The singularity appearing at the wet/dry contact between water and soil surfaces raises difficulties for many numerical methods, but a huge advantage of this problem is that it can be exactly solved [16, 28, 29, 30] and thus, one can make use of its analytic solution to theoretically test the developed numerical method.
Thacker Problem is basically formulated as the PDE system (1) modelling the flow dynamics over a paraboloid type surface

### v2

@@ Abstract
The paper presents an open source software for numerical integration of an extended Saint-Venant model used as a mathematical tool to simulate the water flow from laboratory up to large-scale spatial domains applying physically-based principles of fluid mechanics. Many in-situ observations have shown that vegetation plays a key role in controlling the hydrological flux at catchment scale. In case of heavy rains, the infiltration and interception processes cease quickly, the remaining rainfall gives rise to the Hortonian overland flow and the flash flood is thus initiated. In this context, we also address the following problem: how do the gradient of soil surface and the vegetation influence the water dynamics in the Hortonian flow? The mathematical model and ASTERIX were kept as simple as possible in order to be accessible to a wide range of stakeholders interested in understanding the complex processes behind the water flow on hillslopes covered by plants.
@@
Software Requirements: 32 or 64 bit Linux operating system, gcc compiler, SDL libraries.
Cost: free under GNU license
@@
Availability: http://www.ima.ro/software/ASTERIX_flow_2D.tar.xz
@@ 1 Introduction
Given all that, it is mandatory for all of us to understand the hydrological cycles to properly manage the water resources in order to protect our life and material well-being.
The water distribution in a hydrographic basin is strongly influenced by many factors related to land cover, land use, soil type or soil surface gradients, [4, 5]. Some of these factors play a key role in controlling the destructive effects of the water-related natural hazards. Mathematical modelling of the hydrological processes is an effective way to estimate the risks associated to these hazards. By modelling, one can make scenarios, can evaluate the level of the risk exposure, and can provide important information regarding an economical water use (e.g. to agronomists).
@@
In addition, there are many software products that implement mathematical models, some of them are commercial and some others are free. From the last category, we can mention [11, 12, 13] for hydrology-related R and Python tools and resources. Among the most known models in the class of physical models, we mention SHE [14], MIKE-SHE [15], KINEROS [16], VIC [17], PRMS [18], SWASHES [19].
@@
The novelty of adding a pointwise distributed porosity function to model the presence of plant cover allows the study of water dynamics for various scenarios with different vegetation distributions.
@@
The numerical scheme of the model behind this software was subjected to several internal and external validation tests. Tests such as steady flow over a bump, the oscillating solution of Thacker’s Problem, and the solution to the Riemann Problem of the shallow-water equations were used for the internal validation which consists in comparing the numerical results with the exact solutions of the model. For the external validation, we performed tests to compare the numerical results with laboratory measured data [21, 22]: dam break flow over a bump and in a L-shaped channel, downslope flow through rigid vegetation. A good agreement between the numerical and the exact/measured data was obtained in all these cases. A qualitative analysis emphasizing the similarity between the flow evolution on vegetated soil given by our numerical model and the behavior of the observed phenomenon was also performed to strengthen to confidence in the model.
@@
This paper is organized as follows: in Section 2 we present the PDE system of the Saint-Venant model with vegetation and the numerical scheme built to approximate this model and implemented in ASTERIX. The software is described in Section 3 and some numerical applications add new examples for additional validation on the method and the software used in Section 4. The last section presents final remarks and conclusions.
@@ 2.1 Mathematical model
The Saint-Venant model with porosity we consider in this paper is based on mass and momentum balance equations and takes into account the variation of the soil surface, the plant cover, and the rain as water source. It reads as
$\begin{split}\partial_{t}(\theta h)+\partial_{1}(\theta hv_{1})+\partial_{2}(\theta hv_{2})&=\mathfrak{M},\\ \partial_{t}(\theta hv_{a})+\partial_{1}(\theta hv_{a}v_{1})+\partial_{2}(\theta hv_{a}v_{2})+\theta h\partial_{a}{w}&=\mathfrak{t}^{p}_{a}+\mathfrak{t}^{s}_{a},\quad a=1,2,\end{split}$ (1)
@@
Remark 1.
The Saint-Venant model (1) with porosity addresses the water flow on vegetated surfaces. The structure of the plant cover allows one to use spatial averaging methods, like in theory of flow transport in porous media, see [27] for example. In our case, the “pore space” is the interspace between the plant stems. Urban runoff is another important domain where water flows through obstacles, in this case the “pore space” being the space between buildings.
Apparently, the two domains are similar, but the structure of the pores is very different. This structural difference imposes different mathematical models, and a model performing well in one case may be inadequate for the other case.
There is a very rich literature devoted to the subject of urban runoff, see [28] for a review of models in this area. We bring also to the reader’s attention the works [29, 30, 31, 32] where one can find very important ideas and facts concerning the urban surface water flow models.
@@ 2.2 Numerical scheme
The numerical scheme introduced in [20] to approximate the solution of model (1) casts into the general class of method of lines. We firstly construct a discretization of the spatial variables and differential operators only, and leave the time variable continuous. Then, we define a time discretization scheme to integrate the ODE system previously obtained. A first-order FVM is used for the spatial discretization and a fractional step method for the time discretization [33, 34].
@@
$\begin{array}[]{l}(v_{a})_{(i,j)}:=\displaystyle\frac{v_{a\,i}+v_{a\,j}}{2},\quad a=1,2,\\ (v_{n})_{(i,j)}:=\boldsymbol{v}_{(i,j)}\cdot\boldsymbol{n}_{(i,j)},\end{array}$ (14)
@@
where $l_{(i,j)}$ is the length of the common interface $(i,j)$ between the cells $\omega_{i}$ and $\omega_{j}$. ${\cal S}_{a\,i}$ refers to the gradient of the free water surface, while ${\cal J}_{a\,i}$ and ${\cal L}_{i}$ refer to the flux of the linear momentum and mass, respectively. Combining the spatial discretization with the time fractional step method gives the full discrete scheme associated to (1) of the form
$\begin{split}\sigma_{i}(\theta_{i}h_{i})^{n+1}=&\,\sigma_{i}(\theta_{i}h_{i})^{n}+\triangle t_{n}{\cal L}_{i}((h,\boldsymbol{v})^{n})+\sigma_{i}\triangle t_{n}\mathfrak{M}_{i}(t^{n+1},h^{n+1}),\\ \sigma_{i}(\theta_{i}h_{i}v_{a\,i})^{n+1}=&\,\sigma_{i}(\theta_{i}h_{i}v_{a\,i})^{n}+\triangle t_{n}\left({\cal J}_{a\,i}((h,\boldsymbol{v})^{n})+{\cal S}_{a\,i}((h,w)^{n})\right)-\triangle t_{n}\sigma_{i}{\cal K}_{i}(h^{n+1})|\boldsymbol{v}_{i}^{n+1}|v^{n+1}_{a\,i},\quad a=1,2,\end{split}$ (16)
where a generic $\psi_{i}$ represents the value of $\psi$ on $\omega_{i}$ for $i\in\overline{1,N}$, $\varphi^{n}$ is the value of $\varphi$ at the moment of time $t^{n}$, $\triangle t_{n}:=t^{n+1}-t^{n}$ is the time step, ||$\cdot$|| ⋅ | is the euclidean norm, and
@@
To advance a time step, one must solve the system (16) for each cell $\omega_{i}$, which basically reduces to determining the water depth $h_{i}^{n+1}$ and velocity $\boldsymbol{v}_{i}^{n+1}$ at the time $t^{n+1}$ based on the known values $h_{i}^{n}$, $\boldsymbol{v}_{i}^{n}$, $t^{n}$ and $\triangle t_{n}$. The reader is referred to Appendix A for extended details on solving (16) and for details about the restrictions on the time step $\triangle t_{n}$.
@@ 3 Software description
ASTERIX is a software developed in C on a Linux operating system, being dedicated to the simulation of water flow on vegetated slopes. It is based on the previously described numerical scheme for the mathematical model (1) and uses a regular hexagonal network for the spatial discretization. SDL libraries are required for visualizing the time distribution of water layer depth or velocities. ASTERIX is distributed under GPL free software license: https://www.gnu.org/licenses/gpl-3.0.en.html.
The source code of ASTERIX is a collection of files, most of them containing many specific functions. These component functions can be logically grouped into three categories: Data In, Running and Data Out. One can improve the performance or add new actions by modifying the existent functions or adding new ones. These categories are associated to the three steps of the workflow (see Fig. 3) for operating the software: i) pre-processing, ii) numerical solution, iii) post-processing, respectively. These three steps are relatively independent, depending on the final purposes. The first step is totally independent from the others. The second step depends on the output of the first step, but the simulation can be redone by modifying only some parameters specific to the numerical scheme (e.g maximal time step, final time, etc). The post-processing step depends on the output of the second step. These output data can be also processed by other software. We believe the source code and the accompanying User Guide contain sufficient details to allow a C advised programmer to make the desired modifications.
@@
Domain geometry and the cell partition of the flow domain. The topographic information of the soil provided by a GIS raster is ported to a regular hexagonal raster along this step. One edits the file hexa_parameters.dat in order to establish the number of hexagonal cells on the first row of the hexagonal grid. The number of lines of the hexagonal raster is then automatically calculated depending on the geometry of the GIS raster.
@@
Boundary conditions. The boundary conditions are themselves a rather difficult challenge due to the hyperbolic type of the system of equations. Furthermore, identifying the parts of the border on which different conditions must be established represents another difficulty. The software allows various boundary conditions as free discharge, given mass flux, impermeable wall, and Dirichlet. Depending on the studied problem, ASTERIX allows combinations of such boundary conditions as can be seen on the various tests considered in Section 4. We point out that default boundary condition is free discharge; for other types of boundary conditions, the user must define the configuration of the boundary with the specific boundary conditions for the problem one has to solve. In principle, the software allows spatial and time-varying boundary conditions, but such facilities are not yet implemented at this stage; they must be hard-coded by the user.
@@
Table 1: Folder structure of ASTERIX with the key files
@@
Encoded name in ASTERIX Var & param. Refers to
Unknown water_vector[$\cdot$] $h(t,\boldsymbol{x})$ Water dynamics
@@
External data elevation_vector[$\cdot$] $z(\boldsymbol{x})$ Soil topography
theta_veg[$\cdot$] $\theta(\boldsymbol{x})$ Physics of plant cover
@@
alpha_terr $\alpha_{s}$ Physics of soil
Table 3 presents a list of the main files for running ASTERIX.
Table 3: Main files for running ASTERIX. be - binary executable; bs - Bourne-Again shell script; df - data file
Main Files Type Placement Description
convert be gis_hexa_conversion It converts the data from a DEM file from ’GIS/’ into a hexagonal raster and populates ’ini/” with the input .dat files necessary for the main program ’2D_flow’.
generate be generate_artificial_gis It generates a hexagonal raster for a particular type of artificial relief selected from field_function_parameters.dat and populates ’ini/” with the input .dat files necessary for the main program ’2D_flow’.
2D_flow be main The main executable file for simulating the water flow.
artificial_relief_selection.sh bs main It allows one to select the desired artificial relief from field_function_parameters.dat generates the hexagonal raster through the executable ’generate’.
natural_relief_selection.sh bs main It allows one to select the desired natural relief from ’GIS/’ and generates the hexagonal raster through the executable ’convert’.
vpha.sh bs main It allows one to choose the resolution of the hexagonal network associated by editing the file ’hexa_parameters.dat’ and generates the hexagonal raster through the executable ’generate’ for a specific artificial relief.
vphn.sh bs main It allows one to choose the resolution of the hexagonal network by editing the file ’hexa_parameters.dat’ and converts the data from a DEM file from ’GIS/’ into a hexagonal raster through the executable ’convert’ for a specific natural relief.
vpc.sh bs main It allows one to quickly edit the file ’control_parameters.dat’ and to modify the data from this file.
no_graphics.sh bs main It creates the folder ’main-no_graphics/’ and populates it with the simplified code files (without any graphics) of ASTERIX. The flow simulation with numerical data output only can now be run from this folder.
control_parameters.dat df ini It contains the values of different parameters necessary for running ASTERIX at a given resolution.
hexa_parameters.dat df ini/ini_hexa It contains the values of the parameters for the hexagonal grid resolution.
field_function_parameters.dat df generate_artificial_gis It contains the data for some types of artificial reliefs.
susita.asc df GIS A DEM file with the GIS data for the Susita River basin.
ampoi.asc df GIS A DEM file with the GIS data for the Ampoi’s Valley.
lipaia.asc df GIS A DEM file with the GIS data for the Lipaia’s Valley.
paul.asc df GIS A DEM file with the GIS data for the Paul’s Valley.
Assuming the user has downloaded, decompressed the archive of ASTERIX and compiled the files, we now present a simple example for operating the software on Şuşiţa River basin described in Subsection 4.2.1.
Go to folder ’main/’ and execute the following commands:
\$ ./natural_relief_selection.sh susita
\$ ./2D_flow
The last command runs ASTERIX with a graphic representation of the water depth over time; the user can also run
\$ ./2D_flow -r2
for a graphic representation of velocity field over time. More details on operating the software can be found in the accompanying ’User Guide’.
4 Performance Tests
A suite of examples is presented in [20] for the purpose of validating the numerical scheme. In what follows, we add some new examples for testing the performance of the software. The numerical results are compared with the analytic solutions for some of them and with the laboratory measurements for others. We analyze the response of the software to real terrain data and perform a sensitivity analysis with respect to some model as well as some software parameters.
4.1 Theoretical Tests
4.1.1 Flow on radial symmetric surfaces
@@
The numerical solution was obtained with ASTERIX on a hexagonal raster with approximately $900000$ cells of radius $0.115$ m. A comparison between the analytic and numerical water depth and velocity distributions of the stationary solution of the water flow along a radial section is presented in Fig. 4.
@@
For the errors between the two solutions we used the formulas
$\epsilon_{h}=\frac{1}{N}\sum_{i=1}^{N}{\left|\frac{h_{n}(r_{i})-h_{a}(r_{i})}{h_{a}(r_{i})}\right|},\quad\quad\epsilon_{v}=\frac{1}{N}\sum_{i=1}^{N}{\left|\frac{v_{n}(r_{i})-v_{a}(r_{i})}{v_{a}(r_{i})}\right|},$
Crater type soil Hillock type soil
$\epsilon_{h}$ $0.0055$ $0.0111$
$\epsilon_{v}$ $0.0040$ $0.0018$
4.1.2 Thacker’s Problem
The next 2D experiment, Thacker’s Problem, is a powerful test to check the ability of our software to catch and highlight the propagation of a wet/dry front. The singularity appearing at the wet/dry contact between water and soil surfaces raises difficulties for many numerical methods, but a huge advantage of this problem is that it can be exactly solved [19, 37, 38, 39] and thus, one can make use of its analytic solution to theoretically test the developed numerical method.
Thacker’s Problem is basically formulated as the PDE system (1) modelling the flow dynamics over a paraboloid type surface